OpenStudy (dls):

If the sum of first n terms of a series be 5n^2+2n,then 2nd term=?

hartnn (hartnn):

Arithmetic series ?

OpenStudy (dls):

it says n terms so i guess yes

OpenStudy (anonymous):

S2 - S1 = T2

OpenStudy (dls):

ive forgot how to solve these type of questions,long time..

OpenStudy (anonymous):

Can u Solve nw @DLS

OpenStudy (dls):

no

OpenStudy (anonymous):

By putting n=1 you will get first term, \(S_1\) By n =2 you will get Sum of First two terms : \(S_2\) For Second Term: \[T_2 = S_2 - S_1\]

hartnn (hartnn):

another method.(longer) Sn = n/2(2a+(n-1)d) = n/2(10n+4) so, 2a-d = 4 nd=10n---->d=10 2a = 14---->a=7 T2= 7+d=17

Parth (parthkohli):

If your series is arithmetic, then subtract \(S_1\) from \(S_2\) because \(S_1\) is the first term.

OpenStudy (anonymous):

It is Arithmetic Series..

OpenStudy (shubhamsrg):

one obvious way is as @Yahoo! said,, other long method might be to find what series adds up to this..i.e. we have 5n^2 + 2n => 5(n^2 + n) - 3n =>10n(n+1)/2 - 3n this represents summation(10n - 3) i.e. nth term = 10n-3 so 2nd term = 10*2 -3 = 17 voila! :D

Parth (parthkohli):

Let me explain. Take \(k_1\) to be the first term and \(k_2\) the second. You are doing something like this:\[S_2 - S_1 = k_1 + k_2 - k_1 = k_2\]

hartnn (hartnn):

so many methods!

OpenStudy (dls):

@hartnn @shubhamsrg what is "a" in sn=n/2(2a+(n-1)d)?

hartnn (hartnn):

a= 1st term

OpenStudy (anonymous):

@ParthKohli yes it is..

OpenStudy (dls):

i mean what is the value of a:/

Parth (parthkohli):

\(a\) is the first term.

OpenStudy (dls):

and the value of a =?

OpenStudy (anonymous):

Put n = 1 there you will get a..

OpenStudy (dls):

okay

Parth (parthkohli):

As I told you earlier, \(S_1 = T_1 = a\).

OpenStudy (dls):

okay,thhanks!

Parth (parthkohli):

You're welcome =)